I am stuck on how to solve by substitution. Can someone please help me?
10x+2y=7
y=-5x+3

Answer by ankor@dixie-net.com(22740)

I am stuck on how to solve by substitution. Can someone please help me?
:
10x + 2y = 7
y = -5x + 3
:
Substitute (-5x + 3) for y in the 1st equation, find x:
10x + 2(-5x+3) = 7
:
Multiply what's inside the brackets and you have:
10x - 10x + 6 = 7
:
You can see from this that, this problem can't be solved
0x + 6 = 7, makes no sense
:
Convert the 1st equation to the "y=" form:
2y = -10x + 7
y = -5x + 3.5
;
You can see the lines of the two equations have the same slope so they are
parallel. Never intersect, no solution